OpenStudy (anonymous):

Solve equation: (d+2)(d-5)=-6

OpenStudy (anonymous):

Can anyone help!!!

OpenStudy (anonymous):

i think its d^2 - 5d + 2d - 10 = -6 d^2-3d-10=-6 d^2-3d=4

OpenStudy (anonymous):

and d = -2 or d= 5

OpenStudy (sandra):

well, these are easier to solve if you expand them and then factor them back out once you have something that equals zero

OpenStudy (sandra):

so for example, if you use foil, you get:

OpenStudy (sandra):

d^2 -5d + 2d -10 = -6

OpenStudy (sandra):

d^2 -3d -4 = 0

OpenStudy (sandra):

now that you have the right side equal to zero, you can factor back out

OpenStudy (sandra):

and get (d+1)(d-4) = 0

OpenStudy (sandra):

so once you have a factored left hand side that equals zero, you know that the solutions are the two values that make either of the left hand components zero

OpenStudy (sandra):

since zero times anything is zero

OpenStudy (sandra):

so one of the solutions is d=-1

OpenStudy (sandra):

can you get the other now?